OpenStudy (yumeybitaa):

Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A Launch Area:___(1, 2)___ Point A:___(0, 3)___

OpenStudy (michele_laino):

please you have to apply the subsequent formula: \[\frac{ x-x _{1} }{ x _{2}-x _{1} }=\frac{ y-y _{1} }{ y _{2}-y _{1} }\] where (x_1,y_1) is launch Area, and (x_2,y_2) is Point A, for example

OpenStudy (michele_laino):

@Yumeybitaa please substitute coordinates of your points into above formula and you will get your answer

OpenStudy (anonymous):

Find the slope, Put it into point slope form, solve and put into standard form. I have the correct answer because I did this exact question in virtual school. X+Y=3

OpenStudy (anonymous):

y2-y1/x2-x1 3-2/0-1 1/-1 -1=slope y-y1=m(x-x1) y-2=-1(x-1) y-2=-1x+1 +2 +2 y=-x+3 +x +x y+x=3

OpenStudy (yamyam70):

Good job!

OpenStudy (yumeybitaa):

thank you

OpenStudy (michele_laino):

@Yumeybitaa thank you!